What is the difference of twice the number and 16 is equal to the quotient of three times the number and 8?
Answer by RAY100(1637)

2n-16 = 3n/8,,,,,,multiply thru by 8
.
16n-128 =3n
.
13n = 128
.
n= 9.846
.
check,,2(9.846) -16 = 3(9.846)/8 = 3.692,,,,ok
,
,
If per chance,,what they really meant was 2n-6= 8/3n
.
Multiply by 3n,,,,,6n^2 -18n =8,,,,,6n^2 -18n -8 = 0
.
2(3n^2 -18n -4) =0
.
using the quadratic formula
.
a=3,,,,b=(-9),,,,c=(-4)
.
n = [ -(-9) +/- sqrt{ (-9)^2 -4 (3) (-4)} ] /2(3)
.
n= [ 9 +/- sqrt{ 81 +48}]/6
.
n= [ 9 +/- 11.36]/6
.
n= 3.39, -.39
.
check by inserting in original equation,,,ok